# Tell, Don’t Show!: Language Guidance Eases Transfer Across Domains in Images and Videos

### v1

@@ 4.6 Analysis and Ablations
Importance of source domain images.
While the majority of our accuracy gains stem from the text guidance, the source domain images providing noise-free supervision are also important in learning strong models on the target domain. From Tab. 5, joint training using source and pseudo-labeled target yields improvements of $1.57\%$ for DomainNet and $0.8\%$ on Ego2Exo benchmarks compared to target-only training. More importantly, training jointly on source and target allows deploying a single joint model across both domains as opposed to domain specific models, greatly optimizing inference costs.
DomainNet (Real $\rightarrow$) Ego2Exo
[…]
Only Target 76.03 67.16 65.01 69.35 21.98 42.64 32.31
Joint Train 77.30 68.25 67.35 70.97 22.75 43.53 33.14
Table 5: Importance of Joint Training Using source domain data improves performance on target images while allowing deploying joint models across multiple domains.
@@ 5 Conclusion and Future Directions
We introduce a novel framework called LaGTran to use readily available text supervision and enhance target performance in unsupervised domain transfer scenarios. We first start with the observation that traditional domain alignment approaches yield limited benefits beyond well-understood domain shifts, followed by insights that language provides a semantically richer medium of transfer with reduced domain gaps. This leads to a language-guided transfer mechanism where we train a text classifier on language descriptions from a source domain and then use its predictions on descriptions from a different target domain as supervision for the corresponding images. Despite being conceptually simple and straightforward, we show the remarkable ability of our method to outperform competitive prior approaches on challenging benchmarks like GeoNet and DomainNet for images and proposed Ego2Exo for videos. Through an emphasis on cost-effective or easily producible text supervision, we open new possibilities for advancing domain transfer in scenarios with limited manual supervision. While LaGTran already yields non-trivial boosts in performance, we identify potential areas of further improvement through additionally bridging the domain gaps in text space using language adaptation techniques (Hung et al., 2023) or devising mechanisms to incorporate both image and language guidance that offer complimentary benefits during transfer.
6 Broader Impact Statement
@@ Appendix B Construction of Ego2Exo benchmark
We curate Ego2Exo dataset from the larger Ego-Exo4D dataset (Grauman et al., 2023). Specifically, the keystep annotations provided along with Ego-Exo4D offer fine-grained action labels for several short video clips, called segments, that are manually trimmed from long procedural videos to focus on the keysteps to recognize. We restrict focus to videos from cooking activity, as they include the largest set of segments and labels capturing a diverse pool of actions. Moreover, to ease the difficult task of predicting very fine-grained action classes from short segments (eg: add coffee beans vs. add coffee grounds), we use the provided label hierarchy and manually remap the original 96 annotated actions labels into 22 labels by merging similar classes into a larger, common class. The final category list is as follows:
@@
  2. 2.
@@
  3. 3.
@@
  4. 4.
@@
  5. 5.
    Construct undressed salad
  6. 6.
    Make milk tea
  7. 7.
    Serve
  8. 8.
    Cook
  9. 9.
    Prepare a skillet
  10. 10.
@@
  11. 11.
    Prepare milk (boiled)
  12. 12.
    Brew coffee (french press)
  13. 13.
@@
  14. 14.
    Make salad
  15. 15.
    Prepare milk (frothed)
  16. 16.
    Prepare salad dressing
  17. 17.
    Boil noodles in boiling water
  18. 18.
    Cook noodles in a skillet
  19. 19.
@@
  20. 20.
    Prepare dressing
  21. 21.
    Make chai tea
  22. 22.
    Check paper recipe
To provide text supervision to our algorithm, we use the atomic action descriptions provided in Ego-Exo4D dataset. These descriptions provide a narrative of the events in the video, presented in free-form text from the perspective of a third-party observer. Unlike keystep labels, which are defined between specific start and end times within a video, these text descriptions are associated with distinct timestamps, or a single point in time within the video. To create correspondence mapping between the keystep segments and text descriptions, we adopt the method followed in EgoVLP (Lin et al., 2022): to generate a text description for a segment, we compile all text descriptions that fall within the timestamps defined by the start and end times of that segment. If multiple timestamps exist, we concatenate the corresponding texts; if no timestamps are available, we include no associated text with the segment. Furthermore, we concatenate the annotations provided by multiple annotators in creating the text description.
Our proposed Ego2Exo consists of 10,102 video segments labeled with actions from one of the 22 keysteps, out of which 4200 (41.9%) segments have valid text descriptions available. We split these video segments into two equal groups classwise, and collect ego-videos from one group and exo-videos from the other to create our adaptation benchmark. The same procedure applied to the validation videos yields 3147 validation segments with both ego and exo views. The json file containing our complete set of videos (referenced from Ego-Exo4D dataset), along with annotations and text descriptions is inluded along with the supplementary file.

### v3

@@ 4.4 Analysis and Ablations
Importance of source domain images.
While the majority of our accuracy gains stem from the text guidance, the source domain images providing noise-free supervision are also important in learning strong models on the target domain. We observed that joint training using source and pseudo-labeled target yields improvements of $1.57\%$ for DomainNet and $0.8\%$ on Ego2Exo benchmarks compared to target-only training. More importantly, training jointly on source and target allows deploying a single joint model across both domains as opposed to domain specific models, greatly optimizing inference costs.
@@ 5 Conclusion and Future Directions
We introduce a novel framework called LaGTran to use readily available text supervision and enhance target performance in unsupervised domain transfer scenarios. We first start with the observation that traditional domain alignment approaches yield limited benefits beyond well-understood domain shifts, followed by insights that language provides a semantically richer medium of transfer with reduced domain gaps. This leads to a language-guided transfer mechanism where we train a text classifier on language descriptions from a source domain and then use its predictions on descriptions from a different target domain as supervision for the corresponding images. Despite being conceptually simple and straightforward, we show the remarkable ability of our method to outperform competitive prior approaches on challenging benchmarks like GeoNet and DomainNet for images and proposed Ego2Exo for videos. Through an emphasis on cost-effective or easily producible text supervision, we open new possibilities for advancing domain transfer in scenarios with limited manual supervision. Although LaGTran achieves state-of-the-art performance across several datasets, it relies on external vision-language models for textual guidance in the absence of metadata, potentially constraining its applicability in scenarios where the pre-trained VLM models fail to offer discriminative text supervision. Additionally, while exhibiting fewer domain discrepancies, there remain non-trivial gaps even within the text modality that may reduce the accuracy of pseudo-labels in the target domain, which can be potentially addressed by additionally incorporating text-adaptation mechanisms.
Acknowledgements
We acknowledge support from NSF and a Google Award for Inclusion Research.
Impact Statement
@@ Appendix B Construction of Ego2Exo benchmark
We curate Ego2Exo dataset from the larger Ego-Exo4D dataset (Grauman et al., 2023). Specifically, the keystep annotations provided along with Ego-Exo4D offer fine-grained action labels for several short video clips, called segments, that are manually trimmed from long procedural videos to focus on the keysteps to recognize. We restrict focus to videos from cooking activity, as they include the largest set of segments and labels capturing a diverse pool of actions. Moreover, to ease the difficult task of predicting very fine-grained action classes from short segments (eg: add coffee beans vs. add coffee grounds), we use the provided label hierarchy and manually remap the original 96 annotated actions labels into 24 labels by merging similar classes into a larger, common class. The final category list is as follows:
@@
  4. 4.
    Add water
  5. 5.
    Make dough
  6. 6.
    Make pasta
  7. 7.
    Make salad
  8. 8.
    Make chai tea
  9. 9.
    Make milk tea
  10. 10.
@@
  11. 11.
    Prepare dressing
  12. 12.
    Prepare a skillet
  13. 13.
    Add spring onions
  14. 14.
    Turn off the stove
  15. 15.
    Check paper recipe
  16. 16.
@@
  17. 17.
    Prepare milk (boiled)
  18. 18.
    Construct undressed salad
  19. 19.
    Cook noodles in a skillet
  20. 20.
@@
  21. 21.
@@
  22. 22.
    Boil noodles in boiling water
  23. 23.
@@
  24. 24.
@@
To provide text supervision to our algorithm, we use the atomic action descriptions provided in Ego-Exo4D dataset. These descriptions provide a narrative of the events in the video, presented in free-form text from the perspective of a third-party observer. Unlike keystep labels, which are defined between specific start and end times within a video, these text descriptions are associated with distinct timestamps, or a single point in time within the video. To create correspondence mapping between the keystep segments and text descriptions, we adopt the method outlined in EgoVLP (Lin et al., 2022) as follows: to generate a text description for a segment, we compile all text descriptions that fall within the timestamps defined by the start and end times of that segment. If multiple timestamps exist, we concatenate the corresponding texts; if no timestamps are available, we include no associated text with the segment. Furthermore, we concatenate the annotations provided by multiple annotators in creating the text description.
Our proposed Ego2Exo consists of video segments labeled with actions from one of the 24 keysteps, with corresponding text descriptions for each segment. We split these video segments into two equal groups classwise, and collect ego-videos from one group and exo-videos from the other to create our adaptation benchmark. The same procedure applied to the validation videos yields 3147 validation segments with both ego and exo views. The json file containing our complete set of videos (referenced from Ego-Exo4D dataset), along with annotations and text descriptions is available along with our publicly released code.